OpenStudy (anonymous):

A rectangular garden is x feet wide. The length of the garden is 3 feet more than twice the width. Write a polynomial that represents the area of the garden in square feet.

OpenStudy (anonymous):

x(2x+3) = \[2x^{2}+3x\]

OpenStudy (anonymous):

Ok... i cant really see what you put on x(2x + 3) is that an exponent?

OpenStudy (anonymous):

yes, x(2x+3) = 2(x squared) + 3x

OpenStudy (anonymous):

ok.. and that is it for that one..

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so what u give me is the answer??

OpenStudy (anonymous):

yes, it is a polynomial expression, which is what the question asked for. You don't have enough information to actually have a numerical answer.

OpenStudy (anonymous):

ok